OpenStudy (anonymous):

Solve, square root of 18 multiplied by the square root of 8

OpenStudy (anonymous):

It's: 18 times 8=144 sqrt 144=12 so the answer is 12

OpenStudy (amoodarya):

18=2*3*3 8=2*2*2 so 8*18 =2^4*3^2 root(8*18)=root(2^4*3^2)=2^2*3=12

OpenStudy (anonymous):

Thanks :D